OpenStudy (deadshot):

Solve for x: 3 over 3 x plus 1 over quantity x plus 4 equals 10 over 7x?

OpenStudy (deadshot):

\[\frac{ 3 }{ 3x }+\frac{ 1 }{ x+4 }=\frac{ 10 }{ 7x }\]